# NeuRes: Learning Proofs of Propositional Satisfiability

### v1

@@ 4.4 Assignment Decoder
To extract satisfying assignments, we condition literal embeddings $E^{L}$ on the decoder hidden state $h_{t}$ and then use a sigmoid-activated MLP $\psi$ to assign a truth value $\hat{\mathcal{A}}(l_{i})$ to a literal $l_{i}$ as shown in Eq. 5.
$\hat{\mathcal{A}}(l_{i})=\sigma(\psi(E^{L}_{i}\mathbin{\|}h_{t}))$
@@ 5.1 Dataset
For our training and testing data, we adopt the same formula generation method as NeuroSAT, namely $\textbf{SR}(n)$ where $n$ is the number of variables in the formula. This method was designed to generate a generalized formula distribution that is not limited to a particular domain of SAT problems. To control our data distributions, we vary the range on the number of Boolean variables involved in each formula. For our training data, we use formulas in $\text{{SR}}(U(10,40))$ where $U(10,40)$ denotes the uniform distribution on integers between 10 and 40 (inclusive). To generate our teacher certificates comprising resolution proofs and truth assignments, we use the BooleForce solver (Biere, 2010) on the formulas generated on the SR distribution.
@@ 5.2 Loss Function
We train our model in a supervised fashion using teacher-forcing on solver certificates. During unsat episodes, teacher actions (clause pairs) are imposed over the whole run. The length of the teacher proof dictates the length of the respective episode, denoted as $T$. Model parameters $\theta$ are trained to maximize the likelihood of teacher choices $y_{t}$ thereby minimizing the resolution loss $\mathcal{L}_{Res}$ shown in Eq 6.
@@
[…]
@@ 5.3 Hyperparameters
Our neural solver has several hyperparameters that influence network size, depth, and loss weighting. Embedding and hidden state vectors have a fixed width of 512. We train our model with a batch size of 1 and the Adam optimizer (Kingma & Ba, 2014) for 50 epochs. We linearly anneal the learning rate from $5\times 10^{-5}$ to zero over the training episodes. This yields better results than using a constant learning rate (cf. Appendix E). We use a time discounting factor $\lambda=0.99$ for the episodic loss. We apply global-norm gradient clipping with a ratio of 0.5 (Pascanu et al., 2012).
@@ 6 Generating Resolution Proofs
Figure 5: Performance of all attention variants for unsat problems with different numbers of variables.
@@ 6.1 Attention Variants
To assess the basic resolution performance of NeuRes, we evaluate each attention variant using both static and dynamic embeddings. For this experiment, we perform 32 rounds of message-passing for each input formula.
Table 1: Performance of all attention variants on $\text{{SR}}(U(10,40))$ unsat test formulas.
Variant Static-Embed Dynamic-Embed
Solved (%) p-Len Solved (%) p-Len
Casc-Attn 13.62 1.8 38.98 1.97
Full-Attn 23.68 1.59 87.19 1.74
Anch-Attn 26.28 1.99 48.63 1.6
As shown in Table 1, dynamic embedding is decisively better for all three attention variants, thereby confirming its conceptual merit. While Anchored-Attention leads over other variants under static embeddings, Full-Attention performs significantly better for dynamic embeddings, albeit at the cost of longer proofs on average. We believe that Anch-Attn’s better performance in the static setting can be explained through the full connectivity of its attention grid (proven in Appendix B).
On a closer look, Full-Attn particularly leads the field on difficult formulas that contain more variables. Figure 5 illustrates the performance of the variants for formulas in different numbers of variables. All variants exhibit a nearly monotonic inverse proportionality between the number of variables and the success rate. However, the decline happens at a much slower rate for dynamic Full-Attn than for other variants.
Since dynamic-embedding Full-Attn is the best-performing configuration over in-distribution test settings, we will demonstrate the remaining evaluation experiments exclusively on this variant.
6.2 Number of Message-Passing Rounds
The number of message-passing rounds is an important parameter for the model performance. We experiment with a fixed-length protocol (16, 32, 64), and an adaptive protocol that performs $|V|+1$ rounds to adapt to problem size. This number corresponds to the maximum formula graph diameter (i.e., $|V|+1$) and, hence, guarantees full exploration of graph connections at the end of the protocol. Performance results shown in Table 2 confirm that the adaptive protocol performs significantly better than fixed-length variants. Also, we see that performing too many rounds is in fact detrimental to overall performance, as the success rate drops sharply when more rounds (64) than the number of variables (10 – 40) are performed.
Table 2: Performance for different numbers of message-passing rounds on $\text{{SR}}(U(10,40))$ unsat test formulas.
#Rounds Solved (%) p-Len
16 82.12 1.83
32 87.19 1.74
64 52.11 1.77
$|V|+1$ 89.43 1.71
This finding is of practical significance because it removes the need for optimizing the protocol length as a hyperparamter for a given dataset distribution, which sets NeuRes apart from related approaches such as NeuroSAT. All subsequent experiments exclusively use this adaptive variant.
Table 3: Bootstrapped training data reduction statistics. Gen-2 and Gen-3 metrics are computed relative to their predecessor-shrunk dataset. Gen-1 and Overall are evaluated relative to the original training dataset. Reduction stats are computed on an $\text{{SR}}(U(10,40))$ training dataset while p-Len and success rate are computed on a test set of the same distribution.
Gen-1 Gen-2 Gen-3 Overall
Reduction Depth (max, avg) $(9,2.53)$ $(8,1.59)$ $(5,1.48)$ $(13,2.55)$
Proofs Reduced (%) 57.49 43.12 37.77 75.84
Proof Reduction (max, avg) (%) (86.11, 28.31) (59.78, 12.99) (58.23, 10.74) (86.11, 31.81)
Total Reduction (%) 13.92 5.54 4.57 22.41
p-Len 1.505 1.367 1.23 1.23
Success Rate (%) 76.87 88.27 92.84 92.84
@@ 6.3 Shortening Teacher Proofs with Bootstrapping
This inspired us to devise a bootstrapped training procedure to capitalize on this feature: We pre-roll each input problem using model actions only, and whenever the model proof is shorter than the teacher’s, it replaces the teacher’s in the dataset. In other words, we maximize the likelihood of the shorter proof. In doing so iteratively, the model progressively becomes its own teacher by exploiting redundancies in the teacher algorithm. This introduces a bias towards the model’s own actions while reducing variance by having a more concise learning signal in form of shorter proofs. To mitigate this bias effect, we train several model generations, each trained on the reduced data pool of its predecessor. Each fresh model gets trained on proofs shortened by a different model, which allows the later generations to both exploit the existing reduction insights without having an encoded bias towards them. In the case of a single generation, the latter effect gets throttled, as the model matures, due to that self-bias, which gets reset in each new generation.
The outcome of this bootstrapped training process is summarized in Table 3, where Gen-1 is the model in the first bootstrapping generation. As we can see, the models success rate initially declines compared to the original variant without proof shortening (cf. Table 1). We believe that the shortening entails the model specializing in a subset of the training formulas at the expense of limiting the model’s generality. While this enhances quality of the proofs (on some formulas), it limits the quantity of proofs, i.e., of solved instances. Nevertheless, we find that the performance loss from bootstrapping is gradually overcome through Gen-2 and Gen-3 until eventually beating the non-bootstrapped baseline in terms of both success rate and optimality by a notable margin. The sharp decline in proof length at test time (p-Len) shows that the models transfers the bootstrapped knowledge to unseen test formulas, as opposed to just overfitting on training formulas. In addition to success rate and p-Len, we inspect the reduction statistics of our bootstrapped generations (first three rows of Table 3). Since Gen-x models perform multiple reduction scans over the training dataset, we add a metric of reduction depth computed as the number of progressive reductions made to a proof. To further quantify this effect, we report the maximum and average reduction ratios of reduced proofs relative to teacher proofs. Finally, we report the total reduction made to the dataset size in terms of total proof steps.
In Appendix D, we have compiled additional statistics (cf. Table 5 and Figure 9) on proof shortening during the training process, as well as an example proof reduced by the bootstrapped NeuRes (Figure 10). We only include a small reduction example (from 20 steps to 10 steps) for space constraints, but we have many more examples of much larger reductions (e.g., from $\sim$800800800800 to $\sim$400400400400 steps).
Table 4: Performance of different full solver modes on $\textbf{SR}(40)$. All models were trained on the $\text{{SR}}(U(10,40))$ distribution.
Solver Sat Unsat Total
Solved (%) #Iterations Solved (%) p-Len Solved (%)
Res-Aided 94.4 8.62 70.8 1.33 82.6
Res-Detached 90.2 6.64 67.0 1.5 78.6
NeuroSAT 70 - - - -
@@ 7 Resolution-Aided SAT Solving
In this section, we evaluate the performance of the full integrated SAT solver trained on a hybrid dataset comprising 8K unsatisfiable formulas (and their resolution proofs) and 8K satisfiable formulas (and their satisfying assignments). For the former, timeout ($4\cdot{|\mathcal{P}_{\text{teacher}}|}$) and optimality (p-Len) are measured similarly to previous experiments. For satisfiable formulas, we set the timeout (maximum #trials) to $2\cdot|V|$. In this section, we verify that coupling resolution with sat assignment decoding boosts the overall success rate, by evaluating NeuRes in two functional modes that differ only in how they process sat formulas:
  • Res-Detached: Decodes sat assignments on the original formula with no resolution guidance.
  • Res-Aided: Decodes sat assignments on formulas augmented by resolution-derived clauses.
Table 4 points to three main takeaways. We see that the Res-Aided variant exhibits a higher sat performance than Res-Detached which only considers sat assignments on original clauses. This confirms the merit of augmenting formulas with resolution derivations in accelerating the process of finding satisfying assignments. We also observe a gain in unsat performance for Res-Aided. A potential explanation for this effect could be that skipping resolution during sat episodes induces a misalignment between the shared representations (e.g., embeddings and decoder hidden states) and the resolution network. Lastly, we compare the NeuRes performance on $\textbf{SR}(40)$ to that of NeuroSAT’s. Note that we can only compare the with NeuroSAT in terms of the percentage of sat problems solved, not merely correctly classified. This is because the former is backed with an easy-to-check SAT assignment while the latter is only a single-bit prediction that would require solving the whole problem to verify – making it far less useful. Moreover, there is no unsat success rate for NeuroSAT as it does not produce certificates for unsatisfiability (only predictions). As Table 4 shows, all three NeuRes variants outperform NeuroSAT, with Res-Aided in particular having a remarkably higher sat success rate when trained on the same $\text{{SR}}(U(10,40))$ distribution. This is doubly promising considering NeuroSAT was trained on millions of formulas while NeuRes was trained on only 16K formulas.
7.1 Generalizing to Larger Problems
@@
For this experiment, instead of performing the maximum-score resolution step, we pool the top 50 steps and perform the one that results in the shortest clause. This was found to bring considerable gains to the success rate on larger formula distributions. In Appendix F, we show the base performance obtained without this optimization.
Figure 6 shows the scalability of NeuRes to larger problems by letting it run for more iterations. Compared to NeuroSAT (cf. Selsam et al. (2019)), NeuRes scores a much higher first-try success rate on all 5 problem distributions, and a higher final success rate on all of them except for $\textbf{SR}(40)$ and $\textbf{SR}(40)$ on which our performance is very close to NeuroSAT’s. Particularly, NeuRes shows an adequate first-try success on the 3 largest problem sizes where NeuroSAT solves zero or near-zero problems on the first try.
@@ 8 Conclusion
In this paper, we have presented NeuRes, the first proof-generating neuro-symbolic SAT solver. Unlike other neural SAT solvers in the literature, NeuRes is sound-by-design; therefore it does not need an auxiliary SAT solver to validate its verdicts or certificates. Structurally, NeuRes is a fusion between Pointer Networks and NeuroSAT that advances the functional capabilities of both of them. On the one hand, it extends Pointer Networks by not only generating solutions composed of input tokens but also combinations of a growing pool thereof. On the other hand, NeuRes far surpasses NeuroSAT in terms of performance and data efficiency owing to its resolution-aided mode of operation and certificate-based training process. Moreover, unlike NeuroSAT, NeuRes is a full solver rather than a mere classifier.
Despite its promising benchmark performance, NeuRes cannot solely outperform highly-engineered industrial solvers, as is the case for all neuro-symbolic methods as standalone tools. Nevertheless, NeuRes has demonstrated a unique potential to advance SAT solving through its ability of proof reduction, as proof size is a major challenge in certifying the results of traditional solvers. NeuRes proof reduction is facilitated by a bootstrapped training procedure that uses teacher proofs as a guide as opposed to a golden standard. This relaxation of supervision, together with NeuRes’ data efficiency, naturally points to future work on unsupervised neuro-symbolic methods.
9 Impact Statement
This paper presents work whose goal is to advance the fields of Machine Learning and SAT Solving. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A NeuroSAT Formula Graph Construction
(a) Literal-to-Clause Phase
(b) Clause-to-Literal Phase
@@
NeuroSAT-style formula graphs have two designated node types: clause nodes connected to the literal nodes corresponding to their constituent literals (Selsam et al., 2019). For example, in Figure 7, the clause contents are as follows: $c_{1}=(a\lor b),\quad c_{2}=(a\lor\bar{b}),\quad c_{3}=(\bar{a}\lor\bar{b})$. Each message-passing round involves two exchange phases: (1) Literal-to-Clause, and (2) Clause-to-Literal (and implicitly Literal-to-Complement). This construction is particularly efficient as it allows the message-passing protocol to cover the entire graph connectivity in at most $|V|+1$ rounds where $V$ is the set of variables in the formula.
@@ Appendix B Clause Connectivity Under Static Embeddings
Appendix C Further unsat Performance Breakdowns
(a) Success Rate vs. #Variables
(b) Success Rate vs. #Clauses
(c) p-Len vs. #Variables
(d) p-Len vs. #Clauses
Figure 8: Performance of all attention variants on $\text{{SR}}(U(10,40))$ unsat test formulas. All models were trained on 8K $\text{{SR}}(U(10,40))$ formulas.
Appendix D Teacher Proof Reduction
@@
Figure 9: The progression of the total reduction made by each generation to their given training dataset. We observe that Gen-1 has the highest slope as it operated on the original unreduced dataset whereas Gen-2 and Gen-3 operated on already-reduced datasets where there is less room for further reductions.
Figure 9 shows the progression of the total reduction made by each generation to their given training dataset. We observe that Gen-1 has the highest slope as it operated on the original unreduced dataset whereas Gen-2 and Gen-3 operated on already-reduced datasets where there is less room for further reductions. Indeed, we already see that effect in miniature for all three progressions as their slope tends to attenuate near the end.
(a) Teacher Proof
(b) NeuRes Proof
@@
Appendix E Constant vs. Linearly Annealed Learning Rate
Here we show the effect of linearly annealing the Adam optimizer base learning rate to zero over the number of training episodes on the validation success rate throughout training. As shown in Figure 11, linear annealing brings considerable gains to the model performance. Consequently, we use a linearly annealed learning rate for all our main experiments.
Figure 11: Validation Success Rate With Different LR Schedules
Appendix F Out-Of-Distribution Performance
Figure 12: sat success rate over iterations using base Res-Aided NeuRes

### v2

@@ 4.4 Assignment Decoder
To extract satisfying assignments, we use a sigmoid-activated MLP $\psi$ on top of the literal embeddings $E^{L}$ to assign a truth value $\hat{\mathcal{A}}(l_{i})$ to a literal $l_{i}$ as shown in Eq. 4.
$\hat{\mathcal{A}}(l_{i})=\sigma(\psi(E^{L}_{i}))$ (4)
@@ 5.1 Dataset
For our training and testing data, we adopt the same formula generation method as NeuroSAT [42], namely $\textbf{SR}(n)$ where $n$ is the number of variables in the formula. This method was designed to generate a generalized formula distribution that is not limited to a particular domain of SAT problems. To control our data distributions, we vary the range on the number of Boolean variables involved in each formula. For our training data, we use formulas in $\text{{SR}}(U(10,40))$ where $U(10,40)$ denotes the uniform distribution on integers between 10 and 40 (inclusive). To generate our teacher certificates comprising resolution proofs and truth assignments, we use the BooleForce solver [7] on the formulas generated on the SR distribution.
@@ 5.2 Loss Function
We train our model in a supervised fashion using teacher-forcing on solver certificates. During unsat episodes, teacher actions (clause pairs) are imposed over the whole run. The length of the teacher proof dictates the length of the respective episode, denoted as $T$. Model parameters $\theta$ maximize the likelihood of teacher choices $y_{t}$ thereby minimizing the resolution loss $\mathcal{L}_{Res}$ shown in Eq 5.
@@
$\begin{split}\mathcal{L}_{sat}=\frac{1}{T}\sum\limits_{t}\left[\frac{\gamma^{(T-t)}}{|V|}{\sum\limits_{v}^{V}\text{BCE}(\hat{\mathcal{A}}(v),\mathcal{A}(v))}\right]\end{split}$ (6)
@@ 5.3 Hyperparameters
NeuRes has several hyperparameters that influence network size, depth, and loss weighting. In the experiments we fix the embedding dimension to 128. We train our models with a batch size of 1 and the Adam optimizer [28] for 50 epochs which took about six days on a single NVIDIA A100 GPU. We linearly anneal the learning rate from $5\times 10^{-5}$ to zero over the training episodes. This empirically yields better results than using a constant learning rate. We use a time discounting factor $\lambda=0.99$ for the episodic loss. We apply global-norm gradient clipping with a ratio of 0.5 [37].
@@ 6 Generating Resolution Proofs
Table 1: Performance of all attention variants on unsat $\text{{SR}}(U(10,40))$ test problems.
Variant Static-Embed Dynamic-Embed
Proven (%) p-Len Proven (%) p-Len
Casc-Attn 14.72 1.87 37.33 1.79
Full-Attn 25.38 1.61 95.2 1.67
Anch-Attn 28.72 2.12 60.5 2.28
@@ 6.1 Attention Variants
To assess the basic resolution performance of NeuRes, we evaluate each attention variant using both static and dynamic embeddings. For this experiment, we perform 32 rounds of message-passing for each input formula. As shown in Table 1, dynamic embedding is decisively better for all three attention variants, thereby confirming its conceptual merit. While anchored attention leads over other variants under static embeddings, full attention performs significantly better for dynamic embeddings, albeit at the cost of longer proofs on average. We believe that Anch-Attn’s better performance in the static setting can be explained through the full connectivity of its attention grid (proven in Appendix B). Since dynamic-embedding Full-Attn is the best-performing configuration over in-distribution test settings, we will demonstrate the remaining evaluation experiments exclusively on this variant.
Table 2: Bootstrapped training data reduction statistics. Reduction statistics are computed on the $\text{{SR}}(U(10,40))$ training set while p-Len and success rate are computed on a test set of the same distribution.
Reduction Depth max: $23$, avg: $6.6$
Proof Reduction (%) max: $86.11$, avg: $33.51$
Proofs Reduced (%) 90.08
Total Reduction (%) 31.85
p-Len 1.15
Success Rate (%) 100.0
@@ 6.2 Shortening Teacher Proofs with Bootstrapping
This inspired us to devise a bootstrapped training procedure to capitalize on this feature: We pre-roll each input problem using model actions only, and whenever the model proof is shorter than the teacher’s, it replaces the teacher’s in the dataset. In other words, we maximize the likelihood of the shorter proof. In doing so iteratively, the model progressively becomes its own teacher by exploiting redundancies in the teacher algorithm.
The outcome of this bootstrapped training process is summarized in Table 2. We find that bootstrapping results in notable gains in terms of both success rate and optimality. The sharp decline in proof length (relatively quantified by p-Len) at test time shows that the models transfers the bootstrapped knowledge to unseen test formulas, as opposed to merely overfitting on training formulas. In addition to success rate and p-Len, we inspect the reduction statistics of our bootstrapped variant (first three rows of Table 2). Since the bootstrapped model performs multiple reduction scans over the training dataset, we add a metric for reduction depth computed as the number of progressive reductions made to a proof. To further quantify this effect, we report the maximum and average reduction ratios of reduced proofs relative to teacher proofs. Finally, we report the total reduction made to the dataset size in terms of total number of proof steps.
In Appendix C, we have compiled additional statistics (cf. Table 4) on proof shortening during the training process, as well as an example proof reduced by the bootstrapped NeuRes (Figure 7). We only include a small reduction example (from 20 steps to 10 steps) for space constraints, but we have many more examples of much larger reductions (e.g., from $\sim$800800800800 to $\sim$400400400400 steps).
@@ 7 Resolution-Aided SAT Solving
In this section, we evaluate the performance of our fully integrated model trained on a hybrid dataset comprising 8K unsatisfiable formulas (and their resolution proofs) and 8K satisfiable formulas (and their satisfying assignments). For the unsatisfiable formulas, timeout ($4\cdot{|\mathcal{P}_{\text{teacher}}|}$) and optimality (p-Len) are measured similarly to previous experiments. For satisfiable formulas, we set the timeout (maximum #trials) to $2\cdot|V|$. Ultimately, this section aims to investigate the effect of incorporating a certificate-driven downstream head on the quality of the learnt representations through its impact on the performance of the complementary task, i.e., proving/predicting satisfiability. We use NeuroSAT as our baseline as it employs the same formula embedding architecture. Since NeuroSAT proves sat but only predicts unsat, we train a classification MLP on top of our trained NeuRes model to further showcase the benefit of our representations on prediction accuracy.
Table 3: Performance of full solver mode tested on $\textbf{SR}(40)$ problems and trained on $\text{{SR}}(U(10,40))$ problems where predicted refers to the satisfiability prediction without certificate.
Model Proven (%) Predicted (%)
SAT UNSAT Total SAT UNSAT Total
NeuRes 96.8 99.6 98.2 84.28 99.2 91.65
NeuroSAT [42] 70 - - 73 96 85
Table 3 confirms this main hypothesis. In essence, this result points to the fact that learning signals obtained from training on unsat certificates largely enhance the ability of the neural network to extract useful information from the input formula. This is doubly promising considering NeuroSAT was trained on millions of formulas while NeuRes was trained on only 16K formulas. Lastly, we find that augmenting sat formulas by resolution derivations results in relative improvements ($\sim 2.3\%$) in success rate even though these derivations are attempting to prove unsatisfiability.
8 Generalizing to Larger Problems
@@
Figure 5 shows the scalability of NeuRes to larger problems by letting it run for more iterations. Compared to NeuroSAT [42], NeuRes scores a much higher first-try success rate on all 5 problem distributions, and a higher final success rate on all of them except for $\textbf{SR}(40)$ on which both models nearly score $100\%$. Particularly, NeuRes shows higher first-try success on the 3 largest problem sizes where NeuroSAT solves zero or near-zero problems on the first try.
@@ 9 Conclusion
In this paper, we introduced a deep learning approach for proving and predicting propositional satisfiability. We proposed an architecture that combines graph neural networks with attention mechanisms to generate resolution proofs of unsatisfiability. Unlike methods that merely predict unsatisfiability, our models provide easily verifiable certificates for their verdicts. We demonstrated that our certificate-based training and resolution-aided mode of operation surpass previous approaches in terms of performance and data efficiency, which we attribute to learning better representation.
Despite its promising benchmark performance, our model cannot solely outperform highly-engineered industrial solvers, as is currently the case for all neural methods as standalone tools. The gap between neural networks and symbolic algorithms is still several breakthroughs wide, and our hope is to bring deep learning methods one concrete step closer to filling this gap. For NeuRes, this step is recognizing the immense value of carefully integrating certificates into the model design and training as opposed to using shallow supervision labels. Last but not least, it is worth noting that even at their present state, neural networks stand great potential to advance traditional solvers by combining them in hybrid solvers that utilize the deep long-range dependencies captured by neural networks along with the exploration speed of symbolic algorithms. Moreover, we demonstrated a unique potential to advance SAT solving through proof reduction, as proof size is a major challenge in certifying the results of traditional solvers. This proof reduction is facilitated by a bootstrapped training procedure that uses teacher proofs as a guide as opposed to a golden standard.
Appendix
@@ Appendix A NeuroSAT Formula Graph Construction
NeuroSAT-style formula graphs have two designated node types: clause nodes connected to the literal nodes corresponding to their constituent literals [42]. For example, in Figure 6, the clause contents are as follows: $c_{1}=(a\lor b),\quad c_{2}=(a\lor\bar{b}),\quad c_{3}=(\bar{a}\lor\bar{b})$. Each message-passing round involves two exchange phases: (1) Literal-to-Clause, and (2) Clause-to-Literal (and implicitly Literal-to-Complement). This construction is particularly efficient as it allows the message-passing protocol to cover the entire graph connectivity in at most $|V|+1$ rounds where $V$ is the set of variables in the formula.
@@ Appendix B Clause Connectivity Under Static Embeddings
Appendix C Teacher Proof Reduction